How much water must be evaporated from 240 gallons of a 3% salt solution to produce a 5% salt solution?


Let me see.

240(0.03) - x = 0.05(240 - x)
Is my set up equation correct?


Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

Let water removed be x gallons
Initial quantity = 240 gallons , 3%
240*0.03 = (240-x)*0.05
7.2= 12-0,05x
4.8 = 0.05x
x =96
96 gallons of water must be evaporated to produce a 5% salt solution.







Answer by ikleyn(52776)

.

Your intention was to write equation for the conservation of the dissolved salt mass
in the mixture before and after evaporation.

Right side of your equation (after evaporation) is written correctly:
it is 0.05(240-x).

Left side for the initial salt mass is written incorrectly.
It should be 0.03*240.

Therefore, your equation, in whole, is wrong.


There is one more moment to discuss. It is about units for salt mass.

In this problem, the mass equation for salt is assumed to be written in kilograms.
But for it, the volume of water should be in liters.

So, the conversion factor gallons --> liters should be introduced.

But after introducing this factor, it will present, as a factor, in both sides
of the equation, and, therefore, can be reduced and omitted.

Then the final answer for the volume of water will be in gallons.


You should keep it in your mind, in order for the entire logical line
of the solution be correct in your head.

-----------------

On similar problems, see the lesson
    - How much water must be evaporated
in this site.